# On the bias of the Hoover index estimator: Results for the gamma distribution

### v1

@@ Abstract
The Hoover index is a widely used measure of inequality with an intuitive interpretation, yet little is known about the finite-sample properties of its empirical estimator. In this paper, we derive a simple expression for the expected value of the Hoover index estimator for general non-negative populations, based on Laplace transform techniques and exponential tilting. This unified framework applies to both continuous and discrete distributions. Explicit bias expressions are obtained for gamma population, showing that the estimator is generally biased in finite samples. Numerical and simulation results illustrate the magnitude of the bias and its dependence on the underlying distribution and sample size.
@@ 1 Introduction
The Hoover index is defined in terms of the mean absolute deviation from the population mean and is therefore closely related to $L^{1}$-type dispersion measures (Cowell,, 2011; Yitzhaki,, 1983). Its simplicity and intuitive appeal have led to applications in diverse fields, including income distribution, regional economics, demography, and environmental studies (Rey and Montouri,, 1999; Haining,, 2003). However, from a statistical perspective, relatively little is known about the properties of its empirical estimator, especially in finite samples and under specific distributional assumptions.
@@
The remainder of the paper is organized as follows. Section 2 reviews the definition of the Hoover index and presents several useful characterizations for continuous and discrete distributions. Section 3 establishes technical results based on exponential tilting that are instrumental for the main analysis. Section 4 derives a simple expression for the expected value of the Hoover index estimator and specializes it to gamma population. Section 5 presents a simulation study illustrating the theoretical findings. Section 6 presents a discussion of the implications of our results and directions for future research. The appendix concludes with the derivation of simple expressions for the bias of the Hoover index estimator for discrete Poisson and geometric distributions.
2 The Hoover index: characterizations and special cases
The Hoover index, alternatively referred to as the Robin Hood index or Schutz index, is a quantitative metric designed to assess income inequality, representing the proportion of total income that would need to be reallocated to achieve uniform income distribution.
Definition 2.1.
The Hoover index (Hoover,, 1936) of a non-negative, non-degenerate random variable $X$ with finite mean $\mathbb{E}[X]=\mu>0$ is defined as
$H={\mathbb{E}|X-\mu|\over 2\mu}.$ (1)
@@ 2 The Hoover index: characterizations and special cases
The Hoover index of a non-negative random variable $X$ can be expressed as
@@
$\lvert y_{1}-y_{2}\rvert=\int_{0}^{\infty}\Big[\mathds{1}_{(y_{1},\infty)}(t)+\mathds{1}_{(y_{2},\infty)}(t)-2\,\mathds{1}_{(y_{1},\infty)}(t)\mathds{1}_{(y_{2},\infty)}(t)\Big]\,{\rm d}t,$ (2)
@@
$\lvert X-\mu\rvert=\int_{0}^{\infty}\Big[\mathds{1}_{(0,t)}(X)+\mathds{1}_{(0,t)}(\mu)-2\,\mathds{1}_{(0,t)}(X)\mathds{1}_{(0,t)}(\mu)\Big]\,{\rm d}t.$
@@
The Hoover index of a discrete random variable $X$ supported on $\{0,1,\ldots\}$ admits the characterization
@@
$\int_{0}^{\mu}F(t^{-})\,\mathrm{d}t=\sum_{k=0}^{\infty}\int_{k}^{k+1}F(\lfloor t\rfloor)\,\mathds{1}_{(0,\mu)}(t)\,\mathrm{d}t.$ (3)
@@
$\int_{k}^{k+1}\mathds{1}_{(0,\mu)}(t)\,\mathrm{d}t=\begin{cases}1,&k+1\leqslant\mu,\\[2.0pt] \mu-k,&k<\mu<k+1,\\[2.0pt] 0,&k\geqslant\mu.\end{cases}$
@@
$H={1\over\lambda}\left[\sum_{k=0}^{\lfloor\lambda\rfloor-1}{\Gamma(\lfloor k\rfloor+1,\lambda)\over\Gamma(\lfloor k\rfloor+1)}+(\lambda-\lfloor\lambda\rfloor){\Gamma(\lfloor\lambda\rfloor+1,\lambda)\over\Gamma(\lfloor\lambda\rfloor+1)}\right].$
@@
Since $F(k)=\Gamma(\lfloor k\rfloor+1,\lambda)/\Gamma(\lfloor k\rfloor+1)$ and $\mu=\lambda$, the result follows directly from Proposition 2.2. ∎
@@ 3 Technical results
If $Y_{1}$ and $Y_{2}$ are two non-negative independent random variables, then, for $x_{1},x_{2}>0$,
@@
$F_{Y_{i,x_{i}}}(t)={\mathbb{E}[\mathds{1}_{(0,t]}(Y_{i})\exp\{-Y_{i}x_{i}\}]\over\mathscr{L}_{Y_{i}}(x_{i})},\quad t>0,\quad i=1,2,$
@@
$=\int_{0}^{\infty}\Big\{\mathbb{E}\left[\mathds{1}_{(0,t)}(Y_{1})\exp\{-Y_{1}x_{1}\}\right]\mathscr{L}_{Y_{2}}(x_{2})+\mathbb{E}\left[\mathds{1}_{(0,t)}(Y_{2})\exp\{-Y_{2}x_{2}\}\right]\mathscr{L}_{Y_{1}}(x_{1})$
$-2\mathbb{E}\left[\mathds{1}_{(0,t)}(Y_{1})\exp\{-Y_{1}x_{1}\}\right]\mathbb{E}\left[\mathds{1}_{(0,t)}(Y_{2})\exp\{-Y_{2}x_{2}\}\right]\Big\}\,{\rm d}t.$
@@ 4 Expectation of the Hoover index estimator
$\widehat{H}=\begin{cases}\displaystyle{1\over 2}\left[\dfrac{\displaystyle\sum_{i=1}^{n}|X_{i}-\overline{X}|}{\displaystyle\sum_{i=1}^{n}X_{i}}\right],&\text{if}\,\displaystyle\sum_{i=1}^{n}X_{i}>0\ \text{a.s.},\\[19.91684pt] 0,&\text{if}\,\displaystyle\sum_{i=1}^{n}X_{i}=0\ \text{a.s.},\end{cases}\quad n\in\mathbb{N},\,n\geqslant 2,$ (4)
@@
Let $X_{1},X_{2},\ldots$ be independent copies of a non-negative random variable $X$. The following holds:
@@
$F_{Y_{1,{x\over n-1}}}(t)=\frac{\mathbb{E}\left[\exp\{-xX\}\mathds{1}_{(0,{t\over n-1}]}(X)\right]}{\mathscr{L}_{X}(x)},\quad F_{Y_{2,x}}(t)=\frac{\mathbb{E}\left[\exp\{-x\sum_{j=2}^{n}X_{j}\}\mathds{1}_{(0,t]}(\sum_{j=2}^{n}X_{j})\right]}{\mathscr{L}_{X}^{n-1}(x)},\quad t>0,$
@@
By using the well-known identity ${z}\int_{0}^{\infty}\exp(-zx){\rm d}x=1,$ with $z=\sum_{i=1}^{n}X_{i}$, we have
$\mathbb{E}\left[\dfrac{\displaystyle\sum_{i=1}^{n}|X_{i}-\overline{X}|}{\displaystyle\sum_{i=1}^{n}X_{i}}\right]$ $=\mathbb{E}\left[\sum_{i=1}^{n}|X_{i}-\overline{X}|\int_{0}^{\infty}\exp\left\{-\left(\displaystyle\sum_{i=1}^{n}X_{i}\right)x\right\}{\rm d}x\right]$
$={1\over n}\sum_{i=1}^{n}\int_{0}^{\infty}\mathbb{E}\left[\,\left|(n-1)X_{i}-\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}X_{j}\right|\exp\left\{-(n-1)X_{i}\left({x\over n-1}\right)\right\}\exp\left\{-\left(\displaystyle\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}X_{j}\right)x\right\}\right]{\rm d}x,$ (5)
@@
$\mathbb{E}\left[\dfrac{\displaystyle\sum_{i=1}^{n}|X_{i}-\overline{X}|}{\displaystyle\sum_{i=1}^{n}X_{i}}\right]=\int_{0}^{\infty}\mathbb{E}\left[\,\left|(n-1)X_{1}-\sum_{j=2}^{n}X_{j}\right|\exp\left\{-(n-1)X_{1}\left({x\over n-1}\right)\right\}\exp\left\{-\left(\sum_{j=2}^{n}X_{j}\right)x\right\}\right]{\rm d}x.$ (6)
@@
Finally, by combining (6), (7) with the definition (4) of estimator $\widehat{H}$, the proof follows. ∎
@@
If $X_{1},X_{2},\ldots$ are independent copies of a non-negative discrete random variable $X$, then
@@
[…]
@@ 5 Illustrative simulation study
$\widehat{H}=\begin{cases}\displaystyle\dfrac{1}{2}\left[\dfrac{\displaystyle\sum_{k=1}^{n}|X_{k}-\overline{X}|}{\displaystyle\sum_{k=1}^{n}X_{k}}\right],&\text{if}\ \displaystyle\sum_{k=1}^{n}X_{k}>0\ \text{a.s.},\\[19.91684pt] 0,&\text{if}\ \displaystyle\sum_{k=1}^{n}X_{k}=0\ \text{a.s.},\end{cases}$
@@ 6 Concluding remarks
This paper investigated the finite-sample bias of the Hoover index estimator. By exploiting integral representations based on Laplace transforms and exponential tilting, we derived a general simple expression for its expected value under non-negative populations. Explicit bias formulas were obtained for gamma distribution, showing that the estimator is generally biased in finite samples and that the bias depends on both the sample size and the underlying distribution. The results provide a theoretical basis for numerical bias assessment and suggest the need for bias correction when the Hoover index is used in empirical applications.
@@ A.1 Biased Hoover estimator in Poisson populations
$-{1\over\lambda}\left[\sum_{k=0}^{\lfloor\lambda\rfloor-1}{\Gamma(\lfloor k\rfloor+1,\lambda)\over\Gamma(\lfloor k\rfloor+1)}+(\lambda-\lfloor\lambda\rfloor){\Gamma(\lfloor\lambda\rfloor+1,\lambda)\over\Gamma(\lfloor\lambda\rfloor+1)}\right].$
@@
${1\over\lambda}\left[\sum_{k=0}^{\lfloor\lambda\rfloor-1}{\Gamma(\lfloor k\rfloor+1,\lambda)\over\Gamma(\lfloor k\rfloor+1)}+(\lambda-\lfloor\lambda\rfloor){\Gamma(\lfloor\lambda\rfloor+1,\lambda)\over\Gamma(\lfloor\lambda\rfloor+1)}\right]\\[5.69046pt] \leqslant\text{Bias}(\widehat{H},H)\leqslant\left(1-{1\over n}\right)[1-\exp\{-n\lambda\}]-{1\over\lambda}\left[\sum_{k=0}^{\lfloor\lambda\rfloor-1}{\Gamma(\lfloor k\rfloor+1,\lambda)\over\Gamma(\lfloor k\rfloor+1)}+(\lambda-\lfloor\lambda\rfloor){\Gamma(\lfloor\lambda\rfloor+1,\lambda)\over\Gamma(\lfloor\lambda\rfloor+1)}\right].$

### v2

@@ Abstract
The Hoover index is a widely used measure of inequality with an intuitive interpretation, yet little is known about the finite-sample properties of its empirical estimator. In this paper, we derive a simple expression for the expected value of the Hoover index estimator for general non-negative populations, based on Laplace transform techniques and exponential tilting. This unified framework applies to both continuous and discrete distributions. Explicit bias expressions are obtained for gamma populations, showing that the estimator is generally biased in finite samples. Numerical and simulation results illustrate the magnitude of the bias and its dependence on the underlying distribution and sample size.
@@ 1 Introduction
Formally, the Hoover index is defined in terms of the mean absolute deviation from the population mean.
Definition 1.1.
Let $X$ be a non-negative, non-degenerate random variable with finite mean $\mathbb{E}[X]=\mu>0$. The Hoover index (Hoover,, 1936) is defined as
$H=\frac{\mathbb{E}|X-\mu|}{2\mu}.$ (1)
The Hoover index can therefore be interpreted as one half of the relative mean absolute deviation from the mean, representing the fraction of total income that must be reallocated to achieve perfect equality. Geometrically, the Hoover index corresponds to the maximum vertical distance between the Lorenz curve $L(p)$ and the line of perfect equality, while the Gini coefficient is equal to twice the area enclosed by these two curves (see Figure 1). Its formulation places it within the class of $L^{1}$-type dispersion measures and highlights its close relationship with other inequality indices based on pairwise differences and order statistics (Cowell,, 2011; Yitzhaki,, 1983). Its simplicity and intuitive appeal have made the Hoover index widely applicable across diverse fields, including income distribution, regional economics, demography, and environmental studies (Rey and Montouri,, 1999; Haining,, 2003). However, from a statistical perspective, relatively little is known about the properties of its empirical estimator, especially in finite samples and under specific distributional assumptions.
00.20.20.40.40.60.60.80.81100.20.20.40.40.60.60.80.811ppL(p)L(p)Equality lineLorenz curveHoover gapGini area
Figure 1: Lorenz curve illustrating the Hoover index and the Gini coefficient.
@@
The remainder of the paper is organized as follows. Section 2 reviews the definition of the Hoover index and presents several useful characterizations for continuous and discrete distributions. Section 3 establishes technical results based on exponential tilting that are instrumental for the main analysis. Section 4 derives a simple expression for the expected value of the Hoover index estimator and specializes it to gamma populations. Section 5 presents a simulation study illustrating the theoretical findings. Section 6 presents a discussion of the implications of our results and directions for future research. The appendix concludes with the derivation of simple expressions for the bias of the Hoover index estimator for discrete Poisson and geometric distributions.
2 Characterizations and special cases of the Hoover index
This section presents useful characterizations of the Hoover index and derives closed-form expressions for several important special cases, including gamma, Poisson, and geometric distributions.
@@ 2 Characterizations and special cases of the Hoover index
The Hoover index of a non-negative, non-degenerate random variable $X$ with finite mean $\mathbb{E}[X]=\mu>0$ can be expressed as
@@
$\lvert y_{1}-y_{2}\rvert=\int_{0}^{\infty}\Big[\mathds{1}_{[0,t)}(y_{1})+\mathds{1}_{(0,t)}(y_{2})-2\,\mathds{1}_{[0,t)}(y_{1})\mathds{1}_{(0,t)}(y_{2})\Big]\,{\rm d}t,\quad y_{1}\geqslant 0,y_{2}>0.$ (2)
@@
$\lvert X-\mu\rvert=\int_{0}^{\infty}\Big[\mathds{1}_{[0,t)}(X)+\mathds{1}_{(0,t)}(\mu)-2\,\mathds{1}_{[0,t)}(X)\mathds{1}_{(0,t)}(\mu)\Big]\,{\rm d}t.$
@@
Remark 2.2.
Note that Proposition 2.1 does not exclude the case $\mathbb{P}(X=0)\in(0,1)$.
@@
The Hoover index of a discrete random variable $X$ supported on $\{0,1,\ldots\}$ with finite mean $\mathbb{E}[X]=\mu>0$ admits the characterization
@@
$\int_{0}^{\mu}F(t^{-})\,\mathrm{d}t=\sum_{k=0}^{\infty}\int_{k}^{k+1}F(\lfloor t\rfloor)\,\mathds{1}_{(t,\infty)}(\mu)\,\mathrm{d}t.$ (3)
@@
$\int_{k}^{k+1}\mathds{1}_{(t,\infty)}(\mu)\,\mathrm{d}t=\begin{cases}1,&k+1\leqslant\mu,\\[2.0pt] \mu-k,&k<\mu<k+1,\\[2.0pt] 0,&k\geqslant\mu.\end{cases}$
@@
$H={1\over\lambda}\left[\sum_{k=0}^{\lfloor\lambda\rfloor-1}{\Gamma(k+1,\lambda)\over\Gamma(k+1)}+(\lambda-\lfloor\lambda\rfloor){\Gamma(\lfloor\lambda\rfloor+1,\lambda)\over\Gamma(\lfloor\lambda\rfloor+1)}\right].$
@@
Since $F(k)=\Gamma(k+1,\lambda)/\Gamma(k+1)$ and $\mu=\lambda$, the result follows directly from Proposition 2.3. ∎
@@ 3 Technical results
If $Y_{1}$ and $Y_{2}$ are two independent, non-degenerate, nonnegative random variables, then, for $x_{1},x_{2}>0$,
@@
$F_{Y_{i,x_{i}}}(t)={\mathbb{E}[\mathds{1}_{[0,t]}(Y_{i})\exp\{-Y_{i}x_{i}\}]\over\mathscr{L}_{Y_{i}}(x_{i})},\quad t\geqslant 0,\quad i=1,2,$
@@
$=\int_{0}^{\infty}\Big\{\mathbb{E}\left[\mathds{1}_{[0,t)}(Y_{1})\exp\{-Y_{1}x_{1}\}\right]\mathscr{L}_{Y_{2}}(x_{2})+\mathbb{E}\left[\mathds{1}_{[0,t)}(Y_{2})\exp\{-Y_{2}x_{2}\}\right]\mathscr{L}_{Y_{1}}(x_{1})$
$-2\mathbb{E}\left[\mathds{1}_{[0,t)}(Y_{1})\exp\{-Y_{1}x_{1}\}\right]\mathbb{E}\left[\mathds{1}_{[0,t)}(Y_{2})\exp\{-Y_{2}x_{2}\}\right]\Big\}\,{\rm d}t.$
@@ 4 Expectation of the Hoover index estimator
$\widehat{H}=\displaystyle{1\over 2}\left[\dfrac{\displaystyle\sum_{i=1}^{n}|X_{i}-\overline{X}|}{\displaystyle\sum_{i=1}^{n}X_{i}}\right]\mathds{1}_{\left\{\sum_{i=1}^{n}X_{i}>0\right\}},\quad n\in\mathbb{N},\,n\geqslant 2,$ (4)
@@
Let $X_{1},X_{2},\ldots$ be independent copies of a non-negative, non-degenerate random variable $X$ with finite mean $\mathbb{E}[X]=\mu>0$. The following holds:
@@
$F_{Y_{1,{x\over n-1}}}(t)=\frac{\mathbb{E}\left[\exp\{-xX\}\mathds{1}_{[0,{t\over n-1}]}(X)\right]}{\mathscr{L}_{X}(x)},\quad F_{Y_{2,x}}(t)=\frac{\mathbb{E}\left[\exp\{-x\sum_{j=2}^{n}X_{j}\}\mathds{1}_{[0,t]}(\sum_{j=2}^{n}X_{j})\right]}{\mathscr{L}_{X}^{n-1}(x)},\quad t\geqslant 0,$
@@
By using the well-known identity ${z}\int_{0}^{\infty}\exp(-zx){\rm d}x=1,$ $z>0$, with $z=\sum_{i=1}^{n}X_{i}$, we have
$\mathbb{E}\left[\widehat{H}\right$] $=\mathbb{E}\left[\widehat{H}\mathds{1}_{\left\{\sum_{i=1}^{n}X_{i}>0\right\}}\right]={1\over 2}\,\mathbb{E}\left[\sum_{i=1}^{n}|X_{i}-\overline{X}|\int_{0}^{\infty}\exp\left\{-\left(\displaystyle\sum_{i=1}^{n}X_{i}\right)x\right\}{\rm d}x\right]$
$={1\over 2n}\sum_{i=1}^{n}\int_{0}^{\infty}\mathbb{E}\left[\,\left|(n-1)X_{i}-\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}X_{j}\right|\exp\left\{-(n-1)X_{i}\left({x\over n-1}\right)\right\}\exp\left\{-\left(\displaystyle\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}X_{j}\right)x\right\}\right]{\rm d}x,$ (5)
@@
$\mathbb{E}\left[\widehat{H}\right]={1\over 2}\,\int_{0}^{\infty}\mathbb{E}\left[\,\left|(n-1)X_{1}-\sum_{j=2}^{n}X_{j}\right|\exp\left\{-(n-1)X_{1}\left({x\over n-1}\right)\right\}\exp\left\{-\left(\sum_{j=2}^{n}X_{j}\right)x\right\}\right]{\rm d}x.$ (6)
@@
Finally, by combining (6) and (7), the proof follows. ∎
@@
If $X_{1},X_{2},\ldots$ are independent copies of a non-negative discrete random variable $X$ with finite mean $\mathbb{E}[X]=\mu>0$, then
@@
Since the Hoover index estimator $\widehat{H}$ is scale invariant, it is natural that its respective expectation, $\mathbb{E}[\widehat{H}$], in Item (8), does not depend on the rate $\lambda$.
@@ 5 Illustrative simulation study
$\widehat{H}=\displaystyle{1\over 2}\left[\dfrac{\displaystyle\sum_{i=1}^{n}|X_{i}-\overline{X}|}{\displaystyle\sum_{i=1}^{n}X_{i}}\right]\mathds{1}_{\left\{\sum_{i=1}^{n}X_{i}>0\right\}},$
@@ 6 Concluding remarks
This paper investigated the finite-sample bias of the Hoover index estimator. By exploiting integral representations based on Laplace transforms and exponential tilting, we derived a general simple expression for its expected value under non-negative populations. Explicit bias formulas were obtained for gamma distribution, showing that the estimator is generally biased in finite samples and that the bias depends on both the sample size and the underlying distribution. The results provide a theoretical basis for numerical bias assessment and suggest the need for bias correction when the Hoover index is used in empirical applications. An important direction for future research concerns the relationship between the Hoover index and skewed generalizations of the normal distribution. While the present paper focuses on the sample bias of the Hoover index estimator using exponential tilting and Laplace transform techniques, recent developments indicate that inequality measures may benefit from probabilistic representations based on skew-normal and skew-elliptical families; see Azzalini and Dalla Valle, (1996); Branco and Dey, (2001); Azzalini, (2014). In particular, the Gini index has been successfully linked to closed skew-normal and unified skew-elliptical distributions, which has renewed interest in its sampling properties and inferential behavior Crocetta and Loperfido, (2005); Arellano-Valle et al., (2014). A similar connection for the Hoover index appears promising. For instance, when sampling from a bivariate elliptical distribution, the numerator of the Hoover index follows a univariate skew-elliptical distribution, as a direct consequence of results on linear functions of order statistics Loperfido, (2008). In the Gaussian case, the same statistic is expected to follow a closed skew-normal distribution, by arguments analogous to those employed in the study of the Gini index Crocetta and Loperfido, (2005); Domínguez-Molina et al., (2003). Establishing such connections may provide new insights into the distributional properties of the Hoover index, potentially enhancing its applicability and relevance in both theoretical and applied contexts. Although these developments fall outside the scope of the present work, they represent a natural and worthwhile avenue for future investigation.
@@ A.1 Biased Hoover estimator in Poisson populations
$-{1\over\lambda}\left[\sum_{k=0}^{\lfloor\lambda\rfloor-1}{\Gamma(k+1,\lambda)\over\Gamma(k+1)}+(\lambda-\lfloor\lambda\rfloor){\Gamma(\lfloor\lambda\rfloor+1,\lambda)\over\Gamma(\lfloor\lambda\rfloor+1)}\right].$
@@
${1\over\lambda}\left[\sum_{k=0}^{\lfloor\lambda\rfloor-1}{\Gamma(k+1,\lambda)\over\Gamma(k+1)}+(\lambda-\lfloor\lambda\rfloor){\Gamma(\lfloor\lambda\rfloor+1,\lambda)\over\Gamma(\lfloor\lambda\rfloor+1)}\right]\\[5.69046pt] \leqslant\text{Bias}(\widehat{H},H)\leqslant\left(1-{1\over n}\right)[1-\exp\{-n\lambda\}]-{1\over\lambda}\left[\sum_{k=0}^{\lfloor\lambda\rfloor-1}{\Gamma(k+1,\lambda)\over\Gamma(k+1)}+(\lambda-\lfloor\lambda\rfloor){\Gamma(\lfloor\lambda\rfloor+1,\lambda)\over\Gamma(\lfloor\lambda\rfloor+1)}\right].$